OpenStudy (anonymous):

write a rule for the following function x -1,0,1,2 y1,3,9,27

OpenStudy (anonymous):

my guess is \(3\) so a power

OpenStudy (kc_kennylau):

1. realise that y are all powers of 3

OpenStudy (anonymous):

?

OpenStudy (kc_kennylau):

2. realise that \(\large3^x\) gives you \(\Large\begin{array}{}x&0&1&2&3\\y&1&3&9&27\end{array}\) which is similar to that

OpenStudy (kc_kennylau):

3. Shift \(x\) to the left for one unit. That gives you \(\large3^{x+1}\) which is \(\begin{matrix}x&-1&0&1&2\\y&1&3&9&27\end{matrix}\)

OpenStudy (anonymous):

thanks that -1 is what got me

OpenStudy (kc_kennylau):

\[\Large\begin{array}{c|c|c|c|c}x&-1&0&1&2\\\hline x+1&0&1&2&3\\\hline 3^{x+1}&1&3&9&27\end{array}\]